OpenStudy (anonymous):

solve by the linear combination method 5u + 4v = -18 0.04u - 0.12v = 0.16 Please help, I don't understand how to solve this...

OpenStudy (anonymous):

 so what u do in combination method is that u add the two equations together but before u add them u have to make sure that one variable cancels out after u add them and to do that u multiply or divide all the parts one or both equations. so, i am going to cancel the variable u multiply all parts of second equation by -125 to get -5u at the bottom which could be cancelled out so -125(0.04u - 0.12v = 0.16) -5u+15v=-20  v=-2 then substitute the value of v into any of the equations and solve for u u can do that yourself